# Evaluating Impacts of Traffic Regulations in Complex Mobility Systems Using Scenario-Based Simulations

### v1

@@ 6 Conclusions
In this study, we presented a model-driven, simulation-based approach for the ex-ante evaluation of urban traffic regulation policies, capable of capturing both direct impacts, such as traffic conditions, modal shifts, and emissions, and indirect impacts, including social and economic effects related to mobility. The application of the framework to the real-world case study of Bologna demonstrated that the integration of a physical layer, representing the transportation system and traffic flows, with a social layer, capturing individual behaviors and adaptation to policy interventions, enables systematic exploration of “what-if” scenarios and assessment of the potential consequences of alternative regulatory designs prior to their actual implementation. The proposed approach addresses the complexity inherent in socio-technical urban mobility systems, accounting for nonlinear interactions, uncertainty, and dynamic adaptation of user behavior, while supporting transparent and reproducible generation of counterfactual scenarios. In this way, it enables rigorous assessment of traffic policies before their implementation, offering decision makers actionable insights into potential impacts. The evaluation results show that the proposed framework is able to support the analysis and assessment of the impacts of different possible policies and behavioral parameters. In addition, experimenting with extreme scenarios, in particular for what concerns behaviorsal parameters, appears to be a powerful approach to assess the sensitivity to these parameters: this sensitivity analysis is particularly relevant for the evaluation of the potential impacts of mobility policies, given the difficulty to provide precise and accurate models for behavioral parameters.
Several directions could enhance the proposed framework: spatial disaggregation and multi-scale analysis; dynamic behavioral modeling and long-term adaptation; and integration of social equity dimensions. From a spatial perspective, the current model works at a single-zone level and provides aggregated insights for one urban area. A natural extension would involve a multi-scale approach that divides the selected area into finer spatial units. This would enable the evaluation of the same key indicators at a sub-zone level. Such granularity would reveal localized hotspots and spatial patterns that may be hidden by aggregation. This would support more targeted interventions and allow policymakers to design zone-specific measures.

### v2

@@ 6 Conclusions
In this study, we present a model-driven, simulation-based approach for the ex-ante evaluation of urban traffic regulation policies, capable of capturing both direct impacts, such as traffic conditions, modal shifts, and emissions, and indirect impacts, including social and economic effects related to mobility. The application of the framework to the real-world case study of Bologna demonstrates that the integration of a physical layer, representing the transportation system and traffic flows, with a social layer, capturing individual behaviors and adaptation to policy interventions, enables systematic exploration of “what-if” scenarios and assessment of the potential consequences of alternative regulatory designs before their actual implementation. The proposed approach addresses the complexity inherent in socio-technical urban mobility systems, accounting for nonlinear interactions, uncertainty, and dynamic adaptation of user behavior, while supporting transparent and reproducible generation of counterfactual scenarios. In this way, it enables a rigorous assessment of traffic policies before their implementation, offering decision-makers actionable insights into potential impacts. The evaluation results show that the proposed framework effectively supports the analysis and assessment of the impacts of various possible policies and behavioral parameters. In addition, experimenting with extreme scenarios, in particular for what concerns behavioral parameters, appears to be a powerful approach to assess the sensitivity to these parameters: this sensitivity analysis is particularly relevant for the evaluation of the potential impacts of mobility policies, given the difficulty to provide precise and accurate models for behavioral parameters.
Several directions could enhance the proposed framework. From a spatial perspective, the current model works at a single-zone level and provides aggregated insights for one urban area. A natural extension would involve a multi-scale approach that divides the selected area into finer spatial units. This would enable the evaluation of the same key indicators at a sub-zone level. Such granularity would reveal localized hotspots and spatial patterns that may be hidden by aggregation. This would support more targeted interventions and allow policymakers to design zone-specific measures.